You are considering purchasing stock in a small illiquid enterprise. The most recent net income for the firm was $13,567,000, and there are 568,000 shares outstanding. To make your valuation estimate you will calculate the P/E Ratio for a comparable, publically traded company. The shares of that firm are selling for $42.72. Its most recent net income was $25,950,000 and there are 5,550,000 shares outstanding. Based on this comparable P/E Ratio, how much should you be willing to pay for the shares of your firm of interest?  
Answer #1

PE of publically traded company

PE - Price of share / Earnings per share

42.72 / (25950000/5550000) = 42.72 / 4.676 = 9.136

For small illiquid enterprise

EPS - 13567000 / 568000 = 23.886

PE = Price / Earnings

Price = Earnings * PE = 23.886 * (PE of Publically listed: 9.136) = $218.22 (Price willing to be paid)